### v1

# Unveiling the Invisible: Reasoning Complex Occlusions Amodally with AURA

Abstract
Amodal segmentation aims to infer the complete shape of occluded objects, even when the occluded region’s appearance is unavailable. However, current amodal segmentation methods lack the capability to interact with users through text input and struggle to understand or reason about implicit and complex purposes. While methods like LISA integrate multi-modal large language models (LLMs) with segmentation for reasoning tasks, they are limited to predicting only visible object regions and face challenges in handling complex occlusion scenarios. To address these limitations, we propose a novel task named amodal reasoning segmentation, aiming to predict the complete amodal shape of occluded objects while providing answers with elaborations based on user text input. We develop a generalizable dataset generation pipeline and introduce a new dataset focusing on daily life scenarios, encompassing diverse real-world occlusions. Furthermore, we present AURA (Amodal Understanding and Reasoning Assistant), a novel model with advanced global and spatial-level designs specifically tailored to handle complex occlusions. Extensive experiments validate AURA’s effectiveness on the proposed dataset. The code, model, and dataset will be publicly released.
Figure 1: We introduce AURA, a multi-modal approach designed for reasoning the amodal segmentation mask including both visible and occlusion regions based on the user’s question. AURA can deduct the implicit purpose underneath the question and response with the textual answer along with predicted amodal masks for various objects. Best viewed in color.
**footnotetext: Corresponding author.
1 Introduction
With the development of the Segment Anything Model (SAM) series methods [22, 53, 21], predicting the segmentation mask of visible regions has been largely solved. Amodal segmentation aims to predict the entire shape of an object, including both visible and occluded regions, which is naturally challenging since visual cues for the occluded regions are unavailable. Predicting the intact shape of occluded objects is crucial for various real-world applications, such as autonomous driving [40, 45, 13], robotics [3, 6], and image editing [54, 58, 2].
Existing amodal segmentation approaches have covered the fields including semantic [4, 44], instance [25, 66, 34], and panoptic [37] segmentation. Although existing amodal segmentation methods [15, 40, 11, 52] can address the problem of inferring occluded object shapes to some extent, they lack the ability to integrate with large language models (LLMs) [48, 60, 49], especially in performing reasoning and segmentation based on scene context, limiting their ability in interacting with users and understanding the implicit purpose underneath users questions. Meanwhile, although existing reasoning segmentation methods [23, 50] can work well on predicting the visible segmentation, they cannot predict the complete amodal segmentation mask, limiting their ability to reason in scenarios with complex occlusions. As a matter of fact, there is currently a significant gap in the availability of datasets designed for amodal reasoning segmentation, as well as in the development of algorithms specifically tailored to this task.
To address this gap, we propose a novel task named Amodal Reasoning Segmentation (ARS). As shown in Figure 1, the ARS task aims to interact with the user through textual questions, reasoning in the complex scene, and predicting the accurate segmentation masks with textual answers. For this task, we present a highly generalizable data collection pipeline and construct a new dataset, AmodalReasonSeg, which covers daily life scenes (encompassing both indoor and outdoor environments) derived from the challenging COCOA-cls dataset [66, 11]. The AmodalReasonSeg dataset facilitates object reasoning and segmentation for visible and amodal masks across diverse scenes.
Along with the new task and dataset, we introduce a novel model named AURA (Amodal Understanding and Reasoning Assistant). AURA can understand the user’s input questions, reason about the scene in the image, and predict both visible and amodal segmentation masks simultaneously. Given the challenge of predicting the shape of occluded parts, which are invisible in the image, we propose two key designs to enhance occlusion prediction capabilities. Specifically, we introduce the Occlusion Condition Encoder, which can distinguish whether objects are occluded and assess the occlusion degree comprehensively. Besides, the Spatial Occlusion Encoder can provide spatial guidance on whether each region is occluded, ensuring consistency between visible and amodal mask predictions. It is worth noticing that the proposed AmodalReasonSeg dataset and the designed AURA method support reasoning and mask prediction for various objects within a single-turn conversation, enhancing applicability to real-world scenarios. Furthermore, AURA supports multi-round conversations, increasing its versatility.
Extensive experiments on the proposed AmodalReasonSeg dataset show that AURA outperforms existing methods in prediction accuracy, reasoning capability, and generalizability for complex occlusion scenarios.
The contributions of this paper can be summarized as follows: (1) We propose a novel task: amodal reasoning segmentation, along with a generalizable dataset collection pipeline. We construct a new dataset, AmodalReasonSeg, covering a variety of complex scenes from daily life. (2) We introduce AURA, a multi-modal method integrating vision and language for complex reasoning, capable of predicting both visible and amodal segmentation masks accurately. To the best of our knowledge, AURA is the first method explicitly designed for reasoning segmentation of amodal regions. (3) Extensive experiments demonstrate the effectiveness of AURA through both quantitative and qualitative evaluations.
2 Related Work
2.1 Amodal Segmentation
Amodal segmentation [25] aims to predict the complete shape of an occluded object, encompassing both visible and occluded parts. Several methods [25, 40, 11, 27, 35, 34, 29] are developed by extending the methods designed for visible segmentation or object detection [5, 16, 59, 20, 17, 18], while the relative occlusion order information is leveraged by many methods [63, 61, 38, 58]. Since shape prior knowledge is an essential approach for inferring the shape of the occluded region, several methods [52, 28, 26, 12, 10, 47, 7, 31] are designed to learn and utilize it. For instance, VRSP [52] designs a codebook to store and retrieve shape priors for amodal masks. A3D [26] learns to reconstruct 3D shape priors to guide amodal segmentation. C2F-Seg [12] works in a coarse-to-fine manner equipped with a Vector-Quantized codebook [10]. Moreover, various datasets have been proposed for diverse scenarios including daily life [66, 11], indoor [57, 9, 8, 65], autonomous driving [40, 37, 42], industrial [11], shopping [29], and gaming [15]. However, existing amodal segmentation methods and datasets are limited to using images or videos as input, focusing solely on predicting amodal masks without user interaction or reasoning about implicit user purposes. In this paper, we present the first high-quality dataset for the newly proposed amodal reasoning segmentation task and introduce a Vision-Language Model-based approach specifically designed for this task.
Figure 2: Overall architecture of the proposed AURA. (a) Given an input image and the input question from the user, the Vision Backbone extracts the visual features of the input image, and the Multi-Modal LLM equipped with LoRA is utilized for understanding the input image and textual questions simultaneously and responding with textual answers that include the [SEG] tokens indicating the segmentation masks. (b) For each [SEG], the Prompt Encoder takes its embedding of the Multi-Modal LLM and outputs the refined embeddings corresponding to the [SEG]. (c) Finally, the Visible Mask Decoder predicts the visible mask using each [SEG]’s refined embeddings. The Amodal Decoder predicts the amodal mask using the occlusion-aware embedding predicted by the Occlusion Condition Encoder. A Spatial Occlusion Encoder is designed to constrain the spatial occlusion information of the predicted visible and amodal segmentation masks to be accurate. Best viewed in color.
2.2 Large Models for Reasoning Segmentation
With the rapid advancement of Large Language Models (LLMs) [48, 60, 49] such as GPT-4 [1] and LLaMA [48], language-based approaches have demonstrated impressive zero-shot generalization abilities across diverse real-world applications. Additionally, Vision Language Models (VLMs) [32, 33, 19] like BLIP-2 [24] and LLaVA [33] have achieved state-of-the-art performance in comprehensively understanding both visual and textual information. To leverage the reasoning capabilities of VLMs, LISA [23] introduces the task of reasoning segmentation, which predicts visible segmentation masks by reasoning based on user input questions. Building upon LISA, several recent methods, such as Pixel-LM [43], VISA [55], and LLM-Seg [50], have been proposed to enhance different aspects of the reasoning segmentation task, such as multi-object segmentation, video segmentation, and training efficiency, respectively. Despite these advancements, all existing methods are limited to segmenting only the visible region of target objects, leaving segmenting and analyzing the occluded region untouched. Our work aims to bridge this gap by enabling reasoning and segmenting in an amodal manner, specifically designed to handle complex occlusions in real-world scenarios.
3 Methodology
In this section, we first introduce the definition of the proposed amodal reasoning segmentation task in 3.1. Then the overall architecture of the designed method is presented in 3.2. Next, the special designs for learning and handling occlusion are introduced in 3.3. Finally, the loss functions that organize the training objectives are shown in 3.4.
3.1 Task Definition
For the proposed amodal reasoning segmentation (ARS) task, each image $x_{img}$ of the dataset is associated with various questions $x_{text}$, answers $y_{text}$, and segmentations $y_{seg}$. For each pair, $x_{text}$ is a question from the user input that does not directly point to objects in the image. $y_{text}$ is the ground-truth textual answer that responds to the question and includes the description of one or more objects alongside the [SEG] token representing the position of the segmentation mask. $y_{seg}$ represents the ground-truth visible and amodal segmentation masks ($M_{v}$ and $M_{a}$), corresponding to the [SEG] in the answer $y_{text}$.
3.2 Overall Architecture
As shown in Figure 2, the overall architecture of the proposed AURA includes three stages. Specifically, (1) the Vision Backbone takes the input image $I$ as input for extracting its visual feature $F_{I}$, and the Multi-Modal LLM $\mathcal{E}_{MLLM}$ equipped with LoRA is employed to understand the input image $I$ and the user’s input question $x_{text}$, and predict the textual answer $\hat{y}_{text}$ to respond the question. (2) The Prompt Encoder $\mathcal{E}_{P}$ takes the embedding $e_{mllm}$ of each [SEG] from the hidden states of the Multi-Modal LLM and predicts the refined embedding $e_{r}$ for each [SEG]. (3) Finally the $e_{r}$ and $F_{I}$ are fed into the Visible Mask Decoder $\mathcal{D}_{V}$ for predicting the visible segmentation mask $\hat{M}_{v}$ of each [SEG]. Besides, the $e_{r}$ is used by the Occlusion Condition Encoder $\mathcal{E}_{OC},$ which learns to predict the occlusion-aware embedding $e_{oa}$ that is contained within the occlusion condition information. Then the Amodal Mask Decoder $\mathcal{D}_{A}$ uses the $F_{I}$ and $e_{oa}$ for predicting the amodal segmentation mask $\hat{M}_{a}$ of each [SEG]. The Spatial Occlusion Encoder $\mathcal{E}_{SO}$ is designed to constrain the predicted $\hat{M}_{v}$ and $\hat{M}_{a}$ to conform to the spatial occlusion relationship, which implicitly improves the accuracy of the two predicted masks.
The textual prediction of the Multi-Modal LLM is supervised using the $\mathcal{L}_{text}$, which can be denoted as:
$\mathcal{L}_{text}=\mathcal{L}_{CE}(\hat{y}_{text},{y}_{text}),$ (1)
where $\mathcal{L}_{CE}$ is the Cross-Entropy loss [56]. Moreover, for supervising the visible and amodal mask predictions, the $\mathcal{L}_{mask}$ is used as shown in the following:
$\begin{split}\mathcal{L}_{mask}=\mathcal{L}_{CE}(\hat{M}_{v},{M}_{v})+\mathcal{L}_{CE}(\hat{M}_{a},{M}_{a})+\\ \mathcal{L}_{Dice}(\hat{M}_{v},{M}_{v})+\mathcal{L}_{Dice}(\hat{M}_{a},{M}_{a}),\end{split}$ (2)
where $\mathcal{L}_{Dice}$ is the Dice loss [46].
3.3 Occlusion Learning
Figure 3: The designed occlusion encoders. (a) The Occlusion Condition Encoder learns to embed the occlusion condition of each mask and output the occlusion-aware embedding. (b) The Spatial Occlusion Encoder takes the predicted visible and amodal masks as input and predicts the occlusion-aware spatial map to constrain the spatial relationship between the two predicted masks.
Figure 3 illustrates the designed occlusion encoders for handling occlusion learning in AURA. Detailed network structures, inputs and outputs, and losses are presented.
Occlusion Condition Encoder is designed to improve the ability of the refined embedding $e_{r}$ predicted by the Prompt Encoder for being aware of the occlusion condition of the object(s) corresponding to the [SEG]. Specifically, the $e_{r}$ is fed into the Occlusion Condition Encoder $\mathcal{E}_{OC}$ and predicts the occlusion-aware embedding $e_{oa}$ that is distinguishable about the degree of occlusion. The Occlusion Condition Encoder $\mathcal{E}_{OC}$ additionally predicts the occlusion rate $\hat{r}$ of the object, which reflects the occlusion condition comprehensively and is supervised by $\mathcal{L}_{occ-oc}$ with the ground-truth occlusion rate $r$.
Spatial Occlusion Encoder is proposed to constrain the spatial occlusion conditions between the predicted visible masks $\hat{M}_{v}$ and amodal masks $\hat{M}_{a}$, enhancing the accuracy of visible and amodal segmentation prediction. Specifically, the Spatial Occlusion Encoder $\mathcal{E}_{SO}$ takes both $\hat{M}_{v}$ and $\hat{M}_{a}$ as input and predicts the occlusion-aware spatial map $\hat{M}_{sp}$, which is supervised by $\mathcal{L}_{occ-sp}$ with the ground-truth occlusion-aware spatial map ${M}_{sp}$. The ground truth ${M}_{sp}$ is defined by setting regions belonging to the ground-truth visible mask ${M}_{v}$ as 1 and setting regions belonging to the ground-truth amodal mask ${M}_{a}$ but not ${M}_{v}$ as 2, and setting other background regions as 0. In this approach, the constraint between predicted masks $\hat{M}_{v}$ and $\hat{M}_{a}$ is constructed and contributes to accurate segmentation.
For supervising the predictions of the occlusion condition and the spatial occlusion map, the $\mathcal{L}_{occ}$ is used as shown below:
$\mathcal{L}_{occ}=\mathcal{L}_{MSE}(\hat{r},r)+\mathcal{L}_{CE}(\hat{M}_{sp},{M}_{sp}),$ (3)
where $\mathcal{L}_{MSE}$ is the Mean Squared Error loss.
3.4 Loss Functions
The overall loss function $\mathcal{L}$ consists of four parts, supervising the textual answers $\hat{y}_{text}$ predicted by the Multi-Modal LLM, the predicted visible and amodal masks ($\hat{M}_{v}$ and $\hat{M}_{a}$), and the predictions of two designed occlusion encoders including the occlusion rate $\hat{r}$ and the occlusion-aware spatial map $\hat{M}_{sp}$. The overall loss function can be formulated as:
$\mathcal{L}=\mathcal{L}_{text}+\mathcal{L}_{mask}+\mathcal{L}_{occ}.$ (4)
4 AmodalReasonSeg Dataset
Existing amodal segmentation datasets [40, 11, 57] only contain images/videos and ground-truth masks without language annotations, which cannot fulfill the requirement of the amodal reasoning segmentation task. To this end, it is essential to construct a large-scale dataset with high-quality language annotations, including implicit user questions and the corresponding answers.
In this section, we first introduce the basic information of the newly constructed dataset in 4.1. Next, we demonstrate a general dataset generation pipeline that is generable for diverse scenarios and can be applied to generate new datasets efficiently in 4.2.
Figure 4: The designed generalizable pipeline for dataset generation. (a) Each image and the information of objects in it are gathered and fed into the GPT to generate questions and answers. (b) Then human annotators will check the generated data and improve it manually. Cross-check ensures the quality of the generated data.
4.1 Overview of the Dataset
To construct the new dataset for the amodal reasoning segmentation task, the challenging amodal segmentation dataset COCOA-cls [11, 66] is utilized. The COCOA-cls dataset is built upon the famous COCO [30] dataset, which includes daily scenarios with indoor and outdoor scenes.
For benchmarking the proposed amodal reasoning segmentation (ARS) task, we present the AmodalReasonSeg dataset which is built upon the COCOA-cls dataset. The AmodalReasonSeg dataset includes 3,143 images and 35,494 high-quality question-and-answer pairs corresponding to the visible and amodal segmentation masks. The visible and amodal bounding boxes and the category information are included in our dataset. For each image, there are 11.3 pairs of questions and answers on average annotated to cover diverse potential conversations. Overall, there are 2,018 images and 22,856 pairs of questions and answers in the train set and 1,125 images and 12,638 question-and-answer pairs in the test set.
Compared with existing amodal segmentation datasets, the AmodalReasonSeg dataset supports interacting with users by using textual questions as input, and the answers respond to the implicit purpose underneath the questions and alongside the visible and amodal segmentation predictions. Moreover, it is worth noticing that the answer corresponding to each image could refer to various objects, meaning the proposed dataset supports reasoning and predicting different targets in a single conversation, which is essential in real-world applications.
4.2 Dataset Generation
A generalizable pipeline is designed to generate the new dataset, as shown in Figure 4. Since manually annotating the question-and-answer pairs and the visible and amodal segmentation masks is labor-intensive and time-consuming, we present a semi-automatic pipeline inspired by [43] to keep the efficiency of the annotating process and high-quality annotations. Specifically, the images and annotations, including category information and visible and amodal masks, are reserved from the original dataset. For generating the question-and-answer pairs, each image alongside the annotations is organized by a designed prompt template to be understood by a powerful and versatile large Vision-Language model, which is the ChatGPT-4o in our work that could take both image and text as input and output text response.
The prompt template is carefully designed to require the ChatGPT-4o to generate question-and-answer pairs that satisfy the requirements to ensure the generation of high-quality data. The annotation information of objects for this image is collected and fed to ChatGPT-4o. For each image, ten question-and-answer pairs are required to be generated. Different objects are required to be organized and contained in one answer. Besides, the relative occlusion information between different objects is required to be considered when generating the question-and-answer pairs, which makes the text description in the answer helpful for discriminating the occlusion relationship between objects. The existing annotations from the original dataset of all objects in the image are organized and taken as input for the ChatGPT-4o. Finally, three human annotators check the generated questions and verify if they are qualified. For the ones that need to be improved, human annotators will manually design appropriate questions and answers for these images. Cross-check is utilized between different human annotators to ensure the quality of generated data.
5 Experiments
5.1 Experimental Settings
Dataset. All experiments, including our proposed method and compared methods, are trained on the training set of the proposed AmodalReasonSeg dataset, and evaluated on the validation set of this dataset for fairness. No extra dataset besides the AmodalReasonSeg dataset is used.
Model Structure. We employ the LLaVA-7B-v1-1 [33] as the multi-modal LLM, which is equipped with LoRA for efficient fine-tuning. For the Vision Backbone and the Prompt Encoder, the SAM [22] method with the ViT-H version is utilized. The mask decoder of SAM is employed for the Visible Mask Decoder and Amodal Mask Decoder. Unless specifically specified, all experiments of the proposed AURA method are based on the model structure of using the LLaVA-7B-v1-1 [33] as the multi-modal LLM and the ViT-H version of SAM [22].
Implementation Details. We implement our proposed AURA based on Pytorch [39]. The AdamW [36] optimizer is used with the learning rate set to 1e-3. The WarmupDecayLR learning rate adjustment policy is used with 100 iterations for warmup and then linearly decreases the learning rate to 0. Apart from the ablation studies using 13B models as Multi-Modal LLM Backbone that use two NVIDIA A100 GPUs (80GB Memory) for experiments, all of our experiments are conducted on two NVIDIA V100 GPUs (32GB Memory). The batch size for each GPU card is set to 1. The entire training iterations are 5000, and the gradient accumulation technique is used by setting the accumulation step to 4. The entire training time is about 10 hours.
Evaluation Metrics. The gIoU and cIoU metrics are used for evaluating the segmentation performance of amodal and visible masks. The gIoU is calculated by averaging the Intersection-over-Unions (IoUs) of all images, while the cIoU is calculated by the cumulative intersection dividing the cumulative union of all masks.
Method Amodal Visible
gIoU cIoU gIoU cIoU
Amodal Segmentation Methods
VRSP [52] AAAI’21 ${}_{\text{~{}AAAI'21}}$ 15.18 16.55 17.41 19.75
C2F-Seg [12] ICCV’23 ${}_{\text{~{}ICCV'23}}$ 18.62 18.25 19.12 20.31
SD-Amodal [57] CVPR’24 ${}_{\text{~{}CVPR'24}}$ 20.87 20.09 21.52 23.10
MLLM-based Methods
LISA [23] CVPR’24 ${}_{\text{~{}CVPR'24}}$ 23.05 22.87 24.39 26.81
LLM-Seg [50] CVPR’24 ${}_{\text{~{}CVPR'24}}$ 24.62 23.06 25.19 27.94
GSVA [51] CVPR’24 ${}_{\text{~{}CVPR'24}}$ 42.30 42.87 47.01 48.22
OMG-LLaVA [62] NeurIPS’24 ${}_{\text{~{}NeurIPS'24}}$ 43.92 44.21 49.41 49.38
GLaMM [41] CVPR’24 ${}_{\text{~{}CVPR'24}}$ 45.75 46.82 49.93 54.14
PSALM [64] ECCV’24 ${}_{\text{~{}ECCV'24}}$ 46.81 47.03 50.64 55.19
Baseline 43.55 43.92 48.96 48.63
AURA 47.76 47.32 51.31 55.38
[…]
5.2 Amodal Reasoning Segmentation
Compared Methods. To benchmark the performance of amodal reasoning segmentation on the proposed AmodalReasonSeg dataset, state-of-the-art amodal segmentation methods [52, 12, 57] and multimodal large language model (MLLM)-based methods [23, 50, 51, 62, 41, 64] for reasoning segmentation of visible region are used for comparison. All methods are trained and evaluated on the train set and the val set of our proposed AmodalReasonSeg dataset, ensuring fairness. LLaVA-7B-v1-1 is employed for all MLLM-based methods.
The amodal segmentation methods [52, 12, 57] are designed for taking an image as input and predicting amodal segmentation masks for all objects but cannot understand the user’s question, reasoning, and response with textual answers, and segmentations. Besides, MLLM-based methods [23, 50, 51, 62, 41, 64] are designed to segment the visible regions of each object by taking the image and the user’s question as input.
However, these aforementioned methods can only predict either visible or amodal masks by training twice with ground-truth visible or amodal masks, while the proposed AURA can reason and predict the visible and amodal masks simultaneously. Moreover, MLLM-based methods like LISA can only reason for one object to each question, while the proposed AURA could reason and respond with various objects with textual answers describing them and corresponding segmentation predictions. The baseline method is the proposed AURA without using the designed two occlusion learning modules, including the Occlusion Condition Encoder $\mathcal{E}_{OC}$ and the Spatial Occlusion Encoder $\mathcal{E}_{SO}$.
[…]
Qualitative Results. Figure 5 presents the visualization results of AURA in three cases, showing that AURA can accurately understand the occlusion relation between objects in the image, reason the purpose underneath the question, and predict segmentation masks for various objects. In the first case, although part of the orange is completely occluded by the ice cream on top, and the plate below has a messy pattern, which may cause confusion, AURA still can correctly understand the question and predict an intact amodal segmentation mask for the occluded orange. In the second case, the black cat and the person’s black clothing are easily confused, especially in a dark environment. However, AURA can precisely distinguish between the cat and the person’s black clothing and accurately predict the complete amodal mask of the person, even though a large region of the person is occluded by the cat. In the third case, the front side of the toilet is occluded by the cat to a large extent, while AURA can reason the purpose of the question and predict a high-quality amodal segmentation mask for the occluded toilet and an accurate segmentation mask for the cat.
5.3 Ablation Study
We conduct ablation studies for the proposed AURA from various aspects, including the design of occlusion learning modules, the effectiveness of LoRA, the scale of the Multi-Modal LLM Backbone, and the scale of the Vision Backbone.
Index $\mathcal{E}_{OC}$ $\mathcal{E}_{SO}$ Amodal Visible
gIoU cIoU gIoU cIoU
1 43.55 43.92 48.96 48.63
2 44.86 44.07 49.57 49.10
3 46.95 45.60 50.68 53.72
4 47.76 47.32 51.31 55.38
Table 2: Ablation study of designs for occlusion learning. Best in bold, second underlined.
Occlusion Learning. As shown in Table 2, when the two designed Occlusion Condition Encoder $\mathcal{E}_{OC}$ and Spatial Occlusion Encoder $\mathcal{E}_{SO}$ are not used, the performance of AURA drops in a certain degree. When only using $\mathcal{E}_{OC}$, the occlusion conditions of all objects are learned and can help with our model for comprehensively distinguishing whether objects are occluded and the degree of occlusion for these objects. Besides, when only using $\mathcal{E}_{SO}$, the spatial occlusion distribution can be learned and guide our model to be aware of whether each region is occluded or visible, and constrain the consistency between the predicted visible and amodal masks regarding the spatial occlusion condition. Finally, when both $\mathcal{E}_{OC}$ and $\mathcal{E}_{SO}$ are utilized, the best performance is achieved with the learning of global and spatial occlusion conditions simultaneously.
Method Amodal Visible
gIoU cIoU gIoU cIoU
AURA w/o LoRA 20.84 21.03 25.43 27.95
AURA w/ LoRA 47.76 47.32 51.31 55.38
Table 3: Ablation study of the effectiveness of using LoRA in the Multi-Modal LLM for efficient fine-tuning. w/o and w/ denote without and with the LoRA technique, respectively. Best in bold.
Effectiveness of LoRA. We conduct experiments to evaluate the effectiveness of LoRA, as shown in Table 3. The LoRA [14] technique is utilized to be equipped with the Multi-Modal LLM for efficiently fine-tuning new data while keeping the pre-trained weight of LLM frozen. As shown in Table 3, when not using the LoRA technique, the performance of AURA drops to a certain extent. This is because the Multi-Modal LLM is not pre-trained on the AmodalReasonSeg dataset and has no specific knowledge learned to understand the purpose underneath the user’s question and reason for various occluded objects. Therefore, the Multi-Modal LLM cannot predict textual answers correctly to respond to the user’s question, leading to the performance of visible and amodal segmentation not well.
Multi-Modal LLM Backbone Amodal Visible
gIoU cIoU gIoU cIoU
AURA-7B 47.76 47.32 51.31 55.38
AURA-13B 49.41 49.74 54.13 57.97
AURA-Llama2-13B 52.34 52.95 58.42 60.21
Table 4: Ablation study of the scale of the backbone for the Multi-Modal LLM. Best in bold, second underlined.
Vision Backbone Amodal Visible
gIoU cIoU gIoU cIoU
SAM-ViT-B 42.85 42.76 46.92 50.12
SAM-ViT-L 45.23 45.31 49.18 53.71
SAM-ViT-H 47.76 47.32 51.31 55.38
Table 5: Ablation study of the scale of the Vision Backbone. B, L, and H denote Base, Large, and Huge, respectively. Best in bold, second underlined.
Scale of Multi-Modal LLM Backbone. The effect of the scale of the Multi-Modal LLM backbone is evaluated as shown in Table 4. The AURA-7B, AURA-13B, and AURA-Llama2-13B denote the AURA method with different Multi-Modal LLM backbones corresponding to LLaVA-7B, LLaVA-13B, and LLaVA-Llama2-13B models. These models are different variants of the LLaVA model with different scales of parameters (7 billion and 13 billion) and different generations of Llama. All experiments apart from this ablation study are based on the AURA-7B model. We can observe from Table 4 that the AURA-7B model does not perform as well as the other two models since the fewer number of parameters limits its learning ability for fine-tuning on the proposed AmodalReasonSeg dataset. Moreover, using larger models with 13 billion parameters can bring better performance since the learning capabilities of these models are better. Finally, upgrading the Llama large language model from the first to the second generation further enhances segmentation performance by generating higher-quality answers and refined embeddings used for segmentation.
Figure 5: Visualization results of AURA. Three cases are shown from top to bottom. GT denotes the ground truth. The first two cases show the images, ground-truth amodal and visible masks, and the predicted masks of AURA. In the last case, the predicted amodal mask by AURA for the occludee, which is the occluded toilet, is shown replacing the ground-truth visible mask.
Scale of the Vision Backbone. To evaluate the effect of the scale of the Vision Backbone, which is utilized to extract vision features of the input image and plays an essential role in providing informative features to the segmentation decoders, experiments are conducted, and the results are shown in Table 5. Three scales of the Vision Backbone are compared, including the Base, Large, and Huge versions of SAM, corresponding to 93.7M, 312M, and 641M parameters, respectively. We can observe that the model with a larger number of parameters can achieve better segmentation performance and greater learning ability.
6 Conclusion
In this work, we propose a new task named Amodal Reasoning Segmentation, aiming to predict textual answers alongside visible and amodal segmentation masks in response to the user’s input question that requires understanding and complex reasoning in the image. A generalizable dataset generation pipeline is proposed, which can generate annotations with high quality efficiently. A new dataset named AmodalReasonSeg is presented for daily scenarios where occlusions are challenging for amodal segmentation. Extensive textual question-and-answer pairs are generated, and various objects are referred to in the answers. Moreover, we propose AURA to understand the user’s implicit purpose and reason for the answer, which organizes various objects coherently with accurately predicted visible and amodal segmentation masks. Extensive experiments demonstrate the effectiveness of our proposed AURA on reasoning and segmenting the amodal masks in diverse scenarios.